### v1

@@ Abstract
Context:Rocky planets form by the concentration of solid particles in the inner few au regions of planet-forming disks. Their chemical composition reflects the materials in the disk available in the solid phase at the time the planets were forming. Studying the dust before it gets incorporated in planets provides a valuable diagnostic for the material composition.
Aims:We aim to constrain the structure and dust composition of the inner disk of the young Herbig Ae star HD 144432, using an extensive set of infrared interferometric data taken by the Very Large Telescope Interferometer (VLTI), combining PIONIER, GRAVITY, and MATISSE observations.
Methods:We introduced a new physical disk model, TGMdust, to image the interferometric data, and to fit the disk structure and dust composition. We also performed equilibrium condensation calculations with GGchem to assess the hidden diversity of minerals occurring in a planet-forming disk such as HD 144432.
Results:Our best-fit model has three disk zones with ring-like structures at , , and au. Assuming that the dark regions in the disk at au and at au are gaps opened by planets, we estimate the masses of the putative gap-opening planets to be around a Jupiter mass. We find evidence for an optically thin emission () from the inner two disk zones ( au) at m. Our silicate compositional fits confirm radial mineralogy gradients, as for the mass fraction of crystalline silicates we get around in the innermost zone ( au), mostly from enstatite, while only in the outer two zones. To identify the dust component responsible for the infrared continuum emission, we explore two cases for the dust composition, one with a silicate+iron mixture and the other with a silicate+carbon one. We find that the iron-rich model provides a better fit to the spectral energy distribution. Our GGchem calculations also support an iron-rich and carbon-poor dust composition in the warm disk regions ( au, K).
Conclusions:We propose that in the warm inner regions ( au) of typical planet-forming disks, most if not all carbon is in the gas phase, while iron and iron sulfide grains are major constituents of the solid mixture along with forsterite and enstatite. Our analysis demonstrates the need for detailed studies of the dust in inner disks with new mid-infrared instruments such as MATISSE and JWST/MIRI.
Mid-infrared evidence for iron-rich dust in the multi-ringed inner disk of HD 144432 ††thanks: Based on observations collected at the European Southern Observatory under ESO programmes 190.C-0963(D), 190.C-0963(E), 190.C-0963(F), 0100.C-0278(E), 0103.D-0153(C), 0103.C-0347(C), 0103.D-0153(G), 108.225V.003, 108.225V.011, and 108.225V.006.
@@
Spectral analysis can uncover the mineral buildup of the dust in planet-forming disks (van Boekel et al., 2005; Juhász et al., 2010; Olofsson et al., 2010), thus providing clues about the composition of future planets that formed from that material to be compared with the Solar System. Silicate dust grains in the size range up to an m radius are easily detected in the surface layers of passively heated disks, because of the strong vibrational resonances at infrared (IR) wavelengths that cause emission bands and allow a measure of their chemical composition and size (e.g., Molster & Waters, 2003). Observations have indeed shown the presence of m-sized grains in the disk surface that are substantially larger than those found in the interstellar medium, and of crystalline silicates such as olivines and pyroxenes with very low Fe/Mg ratios, consistent with forsterite (MgSiO) and enstatite (MgSiO) (Henning, 2010). These minerals are also abundant in the inner Solar System bodies, as an example, Mg, Si, and O constitute 59% of the mass of Earth (Morgan & Anders, 1980). However, the single most abundant element present in the inner Solar System planets is Fe (e.g., it makes up 32 % of the mass of Earth, Morgan & Anders, 1980), which is mostly concentrated in the cores of the planets. This implies that during the formation of planets, a large reservoir of iron-containing dust was available. Yet, the detection of iron in the interstellar and circumstellar dust proves to be elusive111There is some observational evidence for metallic iron and iron-enriched silicate dust grains in the circumstellar envelopes of evolved stars (e.g., Habing, 1996; Kemper et al., 2002; Marini et al., 2019).. Metallic iron does not show any conspicuous spectral features at optical and IR wavelengths, in contrast to silicates. The same applies for amorphous carbon dust, which is also presumed to be present in the cosmic dust (e.g., Henning & Salama, 1998; Tielens, 2022).
@@ 3.1 Dust opacities
Species |
Chemical | Lattice | References |
@@
Mg-silicate of olivine stoichiometry |
MgSiO | amorphous | 1 |
Forsterite |
MgSiO | crystalline | 2 |
Mg-silicate of pyroxene stoichiometry |
MgSiO | amorphous | 3 |
Enstatite |
MgSiO | crystalline | 4 |
Iron |
Fe | metallic | 5 |
Carbon |
C | amorphous | 6 |
@@
Metallic iron might become oxidized. Iron oxide solids have several resonances in their mid-IR spectra, mostly between and m. In Fig. 15 we compare the observed Spitzer spectrum of HD 144432 in that spectral region with the opacity curves of FeO (wüstite), FeO (hematite), and FeO (magnetite). As we see no signs of the iron oxide spectral features in the Spitzer data, we conclude that iron oxides are not major constituents of the dust in the inner disk of HD 144432.
@@ 4.1 Runs with specfit
Species |
Zone 1 | Zone 2 | Zone 3 |
@@
Comparing the abundances of MgSiO (crystalline forsterite and amorphous olivine together) vs. MgSiO (crystalline enstatite and amorphous pyroxene together), we see that MgSiO is times more abundant in the innermost zone, and times in the second zone. In the outer zone the same ratio is , so there MgSiO appears to be more abundant, although this is a highly uncertain estimate. The elemental abundances of the disk zones are similar; each zone contains roughly Mg, Si, and O by mass. These numbers relate only to the silicates, as the elemental composition of the hidden dust component is not considered at this stage. The abundance ratios of Mg/Si and Si/O are dex and dex, respectively.
@@ 4.4 Optical depth, temperature, and density profiles
The dust surface densities at the inner edges of the first two zones are in the range of g cm. The iron-rich model gives a total dust mass of M in zone 1, and M in zone 2, while with the carbon-rich model we find M in zone 1 and M in zone 2. The model with carbon has a lower surface density and dust mass, compared to the model with iron. This is because the IR opacity of small carbon grains is roughly an order of magnitude above that of iron, so in order to produce the same amount of radiation less carbon is needed than iron, in terms of mass.
@@ 5.4 Comparison with condensation sequences
The radial distribution of forsterite (MgSiO) and enstatite (MgSiO)171717GGchem cannot model the crystalline structure of the condensates. Thus, in this section we use the terms forsterite and enstatite to encompass both amorphous and crystalline forms of the materials..
@@ 5.5 Hot dust at the inner rim
In Sect. 4.4 we noted that our models prefer a temperature at the inner rim that is K above the sublimation temperature ( K) of regular dust grains. From the GGchem calculation we can evaluate what kind of minerals can endure such high temperatures. Calculations with solar and with the HD 144432 abundances show similar results. The first two condensates forming are tungsten at K, and zirconium dioxide (ZrO) at K, but they are not shown in our plots because of their very low concentrations (). At slightly lower temperatures ( K) corundum (AlO) forms, followed by perovskite (CaTiO, not shown because its low abundance) and gehlenite (CaAlSiO) condensing at K. This condensation sequence is very similar to what was found by Woitke et al. (2018) at bar; the only difference is that in our calculation the condensation occurs at lower temperatures due to the lower pressures in our disk model.
The most abundant condensates (corundum and gehlenite) present at K have concentrations () about 30 times less than the major solid state species iron and forsterite. Moreover, the near-IR opacity of corundum ( cmg) is much lower than that of iron ( cmg). Considering the difference in temperature and in the surface area, we estimate that the m flux from corundum grains at K located at au is only of the flux emitted by iron grains at K at au radius. It remains to be determined, whether that apparently weak signal from corundum and other refractory grains could be detected in observational data.
@@ 6 Summary
GGchem predicts the existence of some refractory minerals which can survive in the K temperature range. The most abundant species under those conditions are corundum (AlO) forming at K, and gehlenite (CaAlSiO) condensing at K. We estimate that the m flux emitted by corundum grains at K located at au is only of the flux emitted by iron grains at K located at au radius. It is still unclear whether that weak emission could be detected with current observational techniques.
@@ Appendix F Degeneracy in the spectral decomposition
Species |
Zone 1 | Zone 2 | Zone 3 |

### v2

@@ Abstract
Context. Rocky planets form by the concentration of solid particles in the inner few au regions of planet-forming disks. Their chemical composition reflects the materials in the disk available in the solid phase at the time the planets were forming. Studying the dust before it gets incorporated in planets provides a valuable diagnostic for the material composition.
Aims. We aim to constrain the structure and dust composition of the inner disk of the young Herbig Ae star HD 144432, using an extensive set of infrared interferometric data taken by the Very Large Telescope Interferometer (VLTI), combining PIONIER, GRAVITY, and MATISSE observations.
Methods. We introduced a new physical disk model, TGMdust, to image the interferometric data, and to fit the disk structure and dust composition. We also performed equilibrium condensation calculations with GGchem to assess the hidden diversity of minerals occurring in a planet-forming disk such as HD 144432.
Results. Our best-fit model has three disk zones with ring-like structures at , , and au. Assuming that the dark regions in the disk at au and at au are gaps opened by planets, we estimate the masses of the putative gap-opening planets to be around a Jupiter mass. We find evidence for an optically thin emission () from the inner two disk zones ( au) at m. Our silicate compositional fits confirm radial mineralogy gradients, as for the mass fraction of crystalline silicates we get around in the innermost zone ( au), mostly from enstatite, while only in the outer two zones. To identify the dust component responsible for the infrared continuum emission, we explore two cases for the dust composition, one with a silicate+iron mixture and the other with a silicate+carbon one. We find that the iron-rich model provides a better fit to the spectral energy distribution. Our GGchem calculations also support an iron-rich and carbon-poor dust composition in the warm disk regions ( au, K).
Conclusions. We propose that in the warm inner regions ( au) of typical planet-forming disks, most if not all carbon is in the gas phase, while iron and iron sulfide grains are major constituents of the solid mixture along with forsterite and enstatite. Our analysis demonstrates the need for detailed studies of the dust in inner disks with new mid-infrared instruments such as MATISSE and JWST/MIRI.
@@
Spectral analysis can uncover the mineral buildup of the dust in planet-forming disks (van Boekel et al., 2005; Juhász et al., 2010; Olofsson et al., 2010), thus providing clues about the composition of future planets that formed from that material to be compared with the Solar System. Silicate dust grains in the size range up to an m radius are easily detected in the surface layers of passively heated disks, because of the strong vibrational resonances at infrared (IR) wavelengths that cause emission bands and allow a measure of their chemical composition and size (e.g., Molster & Waters, 2003). Observations have indeed shown the presence of m-sized grains in the disk surface that are substantially larger than those found in the interstellar medium, and of crystalline silicates such as olivines and pyroxenes with very low Fe/Mg ratios, consistent with forsterite (Mg2SiO4) and enstatite (MgSiO3) (Henning, 2010). These minerals are also abundant in the inner Solar System bodies, as an example, Mg, Si, and O constitute 59% of the mass of Earth (Morgan & Anders, 1980). However, the single most abundant element present in the inner Solar System planets is Fe (e.g., it makes up 32 % of the mass of Earth, Morgan & Anders, 1980), which is mostly concentrated in the cores of the planets. This implies that during the formation of planets, a large reservoir of iron-containing dust was available. Yet, the detection of iron in the interstellar and circumstellar dust proves to be elusive111There is some observational evidence for metallic iron and iron-enriched silicate dust grains in the circumstellar envelopes of evolved stars (e.g., Habing, 1996; Kemper et al., 2002; Marini et al., 2019).. Metallic iron does not show any conspicuous spectral features at optical and IR wavelengths, in contrast to silicates. The same applies for amorphous carbon dust, which is also presumed to be present in the cosmic dust (e.g., Henning & Salama, 1998; Tielens, 2022).
@@ 3.1 Dust opacities
| Species | Chemical | Lattice | References |
@@
| Mg-silicate of olivine stoichiometry | Mg2SiO4 | amorphous | 1 |
| Forsterite | Mg2SiO4 | crystalline | 2 |
| Mg-silicate of pyroxene stoichiometry | MgSiO3 | amorphous | 3 |
| Enstatite | MgSiO3 | crystalline | 4 |
| Iron | Fe | metallic | 5 |
| Carbon | C | amorphous | 6 |
@@
Metallic iron might become oxidized. Iron oxide solids have several resonances in their mid-IR spectra, mostly between and m. In Fig. 15 we compare the observed Spitzer spectrum of HD 144432 in that spectral region with the opacity curves of FeO (wüstite), Fe2O3 (hematite), and Fe3O4 (magnetite). As we see no signs of the iron oxide spectral features in the Spitzer data, we conclude that iron oxides are not major constituents of the dust in the inner disk of HD 144432.
@@ 4.1 Runs with specfit
| Species | Zone 1 | Zone 2 | Zone 3 |
@@
Comparing the abundances of Mg2SiO4 (crystalline forsterite and amorphous olivine together) vs. MgSiO3 (crystalline enstatite and amorphous pyroxene together), we see that MgSiO3 is times more abundant in the innermost zone, and times in the second zone. In the outer zone the same ratio is , so there Mg2SiO4 appears to be more abundant, although this is a highly uncertain estimate. The elemental abundances of the disk zones are similar; each zone contains roughly Mg, Si, and O by mass. These numbers relate only to the silicates, as the elemental composition of the hidden dust component is not considered at this stage. The abundance ratios of Mg/Si and Si/O are dex and dex, respectively.
@@ 4.4 Optical depth, temperature, and density profiles
The dust surface densities at the inner edges of the first two zones are in the range of g cm-2. The iron-rich model gives a total dust mass of M in zone 1, and M in zone 2, while with the carbon-rich model we find M in zone 1 and M in zone 2. The model with carbon has a lower surface density and dust mass, compared to the model with iron. This is because the IR opacity of small carbon grains is roughly an order of magnitude above that of iron, so in order to produce the same amount of radiation less carbon is needed than iron, in terms of mass.
@@ 5.4 Comparison with condensation sequences
The radial distribution of forsterite (Mg2SiO4) and enstatite (MgSiO3)171717GGchem cannot model the crystalline structure of the condensates. Thus, in this section we use the terms forsterite and enstatite to encompass both amorphous and crystalline forms of the materials..
@@ 5.5 Hot dust at the inner rim
In Sect. 4.4 we noted that our models prefer a temperature at the inner rim that is K above the sublimation temperature ( K) of regular dust grains. From the GGchem calculation we can evaluate what kind of minerals can endure such high temperatures. Calculations with solar and with the HD 144432 abundances show similar results. The first two condensates forming are tungsten at K, and zirconium dioxide (ZrO2) at K, but they are not shown in our plots because of their very low concentrations (). At slightly lower temperatures ( K) corundum (Al2O3) forms, followed by perovskite (CaTiO3, not shown because its low abundance) and gehlenite (Ca2Al2SiO7) condensing at K. This condensation sequence is very similar to what was found by Woitke et al. (2018) at bar; the only difference is that in our calculation the condensation occurs at lower temperatures due to the lower pressures in our disk model.
The most abundant condensates (corundum and gehlenite) present at K have concentrations () about 30 times less than the major solid state species iron and forsterite. Moreover, the near-IR opacity of corundum ( cm2g-1) is much lower than that of iron ( cm2g-1). Considering the difference in temperature and in the surface area, we estimate that the m flux from corundum grains at K located at au is only of the flux emitted by iron grains at K at au radius. It remains to be determined, whether that apparently weak signal from corundum and other refractory grains could be detected in observational data.
@@ 6 Summary
GGchem predicts the existence of some refractory minerals which can survive in the K temperature range. The most abundant species under those conditions are corundum (Al2O3) forming at K, and gehlenite (Ca2Al2SiO7) condensing at K. We estimate that the m flux emitted by corundum grains at K located at au is only of the flux emitted by iron grains at K located at au radius. It is still unclear whether that weak emission could be detected with current observational techniques.
@@ Appendix F Degeneracy in the spectral decomposition
| Species | Zone 1 | Zone 2 | Zone 3 |